Question 4, A glider flies south east (at 320.0 degrees) at 25.0 Km/h. a wind blows at 12.0 Km/h from 15.0 degrees south of west. What is the new velocity of the glider with respect to the ground in standard position? The Applied Physics book answer is 33.4 Km/h at 337.1 degrees. Please show me the steps to derive at the book answer.
Answer #1

Sum the horizontal components:

25cos320 + 12cos(195-180) = 25cos320 + 12cos15 = 19.7+11.6 = 30.74 km/h

Sum vertical components

25sin320 + 12sin15 = -16.1 +3.1 = -13 km/h

√30.74² + 13²) = 33.4km/h

Direction = 360 + arctan(-13/30.74) =360 -22.9° = 337.1 degrees